Grade 11Physical Chemistry

2gm of CaCO3 on heating will give CO2 at STP equal to?

Manthan
8 Years agoGrade 11

1 Answer

Arun
ApprovedApproved Tutor Answer8 Years ago

CaCO3 -> CaO + CO2
The first thing to do here is to calculate the relative formula mass of CaCO3 and CO2

RFM of CaCO3 = 40 g + 12 g + (16 g x 3) = 100 g

RFM of CO2 = 12 g + (16 g x 2) = 44 g

The decomposition equation shows 1 CaCO3 goes to 1 CaO and 1 CO2.

We have 2 gm of CaCO3

2 g / 100 g = 0.02

In order to get the mass of CO2 we need to times the RFM of CO2 by 0.02 because of the ratio of CaCO3 to CO2:

44 g x 0.02 = 0.88 g

0.88 g of CO2 are obtained from decomposition of 2 g of CaCO3